### v1

# Conditional success of adaptive therapy: the role of treatment-pausing thresholds revealed by mathematical modeling SUMMARY

Conditional success of adaptive therapy: the role of treatment-pausing thresholds revealed by mathematical modeling
SUMMARY
Adaptive therapy (AT) improves cancer treatment by controlling the competition between sensitive and resistant cells via treatment holidays. This study highlights the critical role of treatment-pausing thresholds in AT for tumors composed of drug-sensitive and resistant cells. Using a Lotka-Volterra model, the research compares AT with maximum tolerated dose therapy and intermittent therapy, showing that ATβs success largely depends on the threshold at which treatment is paused and resumed, as well as the competition between sensitive and resistant cells. Three scenarios of comparison between AT and others are identified: uniform-decline, conditional-improve, and uniform-improve, illustrating that optimizing the treatment-pausing threshold is crucial for AT effectiveness. Tumor composition, including initial tumor burden and the proportion of resistant cells, influences outcomes. Adjusting threshold values enables AT to suppress resistant subclones, preserving sensitive cells, ultimately improving progression-free survival. These findings underscore the importance of personalized strategies for cancer management and enhancing long-term therapeutic outcomes.
KEYWORDS
Adaptive therapy, treatment-holiday threshold, cancer dynamics, mathematical modeling, treatment outcomes
INTRODUCTION
Cancer remains one of the leading causes of death worldwide 1, with its biological complexity and heterogeneity posing significant challenges to conventional therapies such as chemotherapy, radiotherapy, and even immunotherapy 2, 3, 4. Therapeutic resistance, a central obstacle in cancer treatment, arises from the therapeutic selection pressure of pre-existing or induced resistant tumor cells 5, 6, 7, 8, 9. This process drives the evolution of tumor tissue, ultimately leading to the dominance of primary or induced refractory subpopulations, and subsequent relapse and treatment failure 10, 11, 12. The inherent heterogeneity and plasticity of tumor cells often render complete eradication exceedingly difficult, highlighting the necessity of strategies that address the entire tumor population and its evolutionary dynamics rather than selectively targeting certain subpopulations4, 13, 14.
Recently, ecological theory has provided an innovative framework for understanding tumor progression, conceptualizing tumors as ecosystems where drug-sensitive and resistant cells compete for resources 15, 16, 17, 18, 19, 20, 21, 22, 23. Within this ecological framework, cancer cells under drug treatment activate regenerative or protective mechanisms to survive near-lethal stress, enabling them to enter a stem cell-like state and transmit resistance to their progeny 24, 25. This adaptation imposes costs, such as reduced proliferation rates, decreased environmental carrying capacity, and intensified interclonal competition 26, 27.
The conventional maximum tolerated dose (MTD) therapy aggressively eliminates sensitive cell populations. However, this strategy inadvertently disrupts competitive equilibria, allowing resistant subclones to proliferate uncontrollably 28, 29, 30, 16, 31. This unintended consequence of MTD therapy suggests the limitations of conventional eradication-focused paradigms and highlights the need for therapeutic strategies that preserve sensitive cells to maintain ecological suppression of resistant populations 15, 16, 32, 33.
Adaptive therapy (AT), a dynamic treatment strategy, addresses this challenge by adjusting therapeutic intensity and scheduling in response to real-time tumor burden 16, 34, 29. Unlike fixed-dose regimens, AT employs a feedback-controlled approach: treatment is paused when tumor burden declines to certain lower threshold and resumed upon regrowth to an upper threshold. By retaining a reservoir of sensitive cells, this intermittent dosing modality suppresses resistant subclone expansion and prolongs progression-free survival, as evidenced by clinical successes in prostate cancer and other malignancies.
Despite significant advancements 31, 26, 35, 36, 37, further research is needed to refine the optimization of treatment pause thresholds, a core parameter governing AT efficacy, as most existing studies focus on the efficiency of adaptive therapy under predefined thresholds. Several studies have explored the effects of varying thresholds on AT outcomes across different scenarios 38, 37, 39, 36, 40, 41, providing valuable insights that underscore the importance of threshold adjustments and establishing a solid foundation for further investigation. However, a deeper understanding of the dynamic mechanisms driving context-dependent outcomes under heterogeneous physiological conditions remains essential. A systematic analysis of how different thresholds will give rise to distinct effects and the underlying dynamical mechanisms, which in turn will enhance conventional empirical threshold selection to achieve an optimal balance between tumor suppression and resistance mitigation. Moreover, it can further facilitate the development of personalized strategies tailored to different conditions. Consequently, quantitative modeling of threshold-driven tumor evolutionary dynamics is essential for advancing precision in adaptive therapy design.
Therefore, this study leverages a mathematical model of tumor evolutionary dynamics to analyze and explain the effects of threshold on the outcomes of AT under various conditions of competitive interactions between sensitive and resistant subpopulations, elucidating the biophysical principles underlying threshold-mediated resistance control. The findings are expected to provide theoretical validation for personalized adaptive therapy protocols, facilitating a paradigm shift toward dynamically optimized cancer management.
In summary, this study, as a state-of-the-art work, systematically investigates the impact of treatment pause thresholds on adaptive therapy outcomes using a LotkaβVolterra model that stimulates the competitive dynamics between drug-sensitive and drug-resistant tumor cell populations. By quantifying the relationship between treatment-pausing threshold and employing time to progression (TTP) as an indicator of PFS, the analysis demonstrates that AT efficacy critically depends on the competition between two subpopulations of cancer cells: while weak competition limits the benefits of AT, strong competition can even allow for the indefinite delay of disease progression. Furthermore, this work also reveals that AT is not uniformly superior to MTD, as its efficiency depends on tumor composition (initial conditions) and patientsβ personal conditions (parameters). By comparing the PFS by AT and MTD, it identifies three distinct scenarios, 1) that the which are that AT uniformly improves the PFS than MTD regardless of thresholds, 2) that AT has only conditionally improved the PFS and 3) that AT has uniformly worsened the PFS. These findings highlight the necessity of precise threshold tuning to optimize treatment outcomes. Collectively, these findings provide a theoretical foundation for refining empirical threshold selection and advancing personalized adaptive therapy protocols, thereby paving the way for dynamically optimized cancer management.
RESULTS
Model description
To investigate the effects of different treatment strategies, the tumor dynamics under therapeutic conditions are modelled by the widely used Lotka-Volterra (LV)-type equations 42, 18, 36, 43, 44, 45, 46 because the cancer cells behave more similar to unicellular organism rather than normal human cells 17, 47, 48, 49. The cancer cells are further assumed to be classified into two types: drug-sensitive cells (), which can be killed by drugs, and drug-resistant cells (), which cannot be eliminated by drugs. These two cell types compete within the cancer microenvironment (Figure 1A), and therefore the LV model is employed as the following ordinary differential equations (ODEs) (1)-(3).
where denotes the dose curve of therapeutic agents and and represent the burdens of the two types of cells respectively. Equation (1) describes the growth and death of sensitive cells, which means that in the absence of therapy, their proliferation is regulated by intrinsic rate , logistic competition and the death rate is . The drug-dependent suppression of sensitive cells is represented by the factor where is the killing strength of therapeutic agents on sensitive cells. Although this factor is multiplied with the proliferation terms, it is the integrated effective term taking both suppression of proliferation and increasing death rate for simplicity. Equation (2) governs resistant cells, which grow logistically at rate and die at constant rate , unaffected by the drug. is the carrying capacity of the cancer microenvironment, indicating the maximum number of cells that can be supported.
Moreover, the model equations (1)β(2) incorporate asymmetric competition between drugβsensitive and drugβresistant cells, characterized by the parameters and . Here, quantifies the inhibitory effect of drugβresistant cells on the growth of sensitive cells, while represents the impact of drugβsensitive cells on the growth of resistant cells. Specifically, a higher value of implies that resistant cells more strongly reduce the growth potential of sensitive cells, possibly due to resource competition or indirect suppression via secreted factors (e.g., growth factors, cytokines). Similarly, a higher value of indicates that sensitive cells exert a stronger influence on the growth of resistant cells, potentially through competition for space, nutrients, or other resources within the tumor microenvironment.
These competitive interactions determine the dominance relationships between the cell populations. For example, in the absence of therapy, coexistence of the two populations is possible under weak competition (i.e., and ), whereas strong competition leads to bistability 36, 18, 50, 51. Specifically, if the resistant cell population is competitively dominant (i.e., and ), the sensitive cell population will eventually become extinct; conversely, if and , the resistant cell population will be eliminated 36. In addition, the competition coefficients also significantly influence treatment outcomes, with a competitive advantage for sensitive cells being crucial for eradicating resistance 52, 53, 52, 54. Thus, it is necessary to take these competitive dynamics into account in order to design better strategies.
Different therapeutic strategies are characterized by distinct profiles of the dosing function . Two widely used therapeutic approaches and two regimens of adaptive therapy are considered in this work, including maximum tolerated dose, intermittent therapy (IT), dose-skipping adaptive therapy (AT-S) and dose-modulating adaptive therapy (AT-M) (Figure 1B). The conventional maximum tolerated dose approach seeks to maximize cancer cell eradication by administering the highest dose that a patient can tolerate based on acceptable toxicity levels, which can be defined by continuous administration as (4).
(MTD) | (4) |
As continuous high-intensity dosage leads to the development of therapeutic resistance through Darwinian selection pressures 29, as well as that high dosage of therapeutic agents causes severe toxicity, the therapy holiday, intermittent therapy regimens incorporating treatment holidays were proposed, which strategically suspends drug administration during predefined recovery periods and emerged as a clinically utilized strategy for balancing therapeutic efficacy with toxicity management 43. The IT strategies can be characterized as deterministic treatment cycles as (5).
(IT) | (5) |
where is the fixed period of one cycle of treatment, is the duration of drug administration and represents the number of cycles.
As described before, AT dynamically adjusts cancer treatment based on tumor response to prolong efficacy and manage resistance. Two primary approaches within adaptive therapy are dose-skipping and dose-modulating regimens. In dose-skipping adaptive therapy, treatment is administered at the maximum tolerated dose until the tumor burden shrinks to a predetermined threshold (), for example, 50% of the baseline burden; therapy is then paused, allowing the tumor to regrow to a specific threshold (), usually, 100% of the baseline, before resuming treatment. In dose-modulating adaptive therapy, treatment doses are adjusted at regular intervals based on tumor response: increasing the dose if the tumor grows and decreasing it if the tumor shrinks. This strategy seeks to tailor therapy dynamically to tumor behavior, potentially reducing toxicity and managing resistance more effectively. Both approaches exploit the competitive dynamics between treatment-sensitive and resistant cancer cells to enhance long-term treatment outcomes and have already succeeded in trials or experiments 35, 29, 55, 56, 31. Based on the above description, the two AT regimens can be formulated as following equations (6) and (7) respectively.
(AT-S) | (6) | |||
(AT-M) | (7) |
where is the initial tumor burden and is the derivative of representing whether the tumor burden is growing or shrinking. The threholds satisfies and . Notably, the dosing strategies (6) and (7) represent state-dependent switching mechanisms rather than closed-form mathematical definitions. This implementation induces history-denpendent therapeutic decision and consequently, the governing equations (1)-(3) constitute delay differential equations with history-dependent terms .
As a state-of-the-art investigation, this study provides comprehensive parametric analysis rather than case-specific parameter estimation. The parameter ranges derive from established experimental and clinical researches 35, 36, 57, 58, 59. The proliferation rates of sensitive and resistant cells, i.e. and , are derived from in vitro experimental data 35, while the drug-induced killing strength is sourced from the estimation based on several expriments of cell lines57. Given the critical clinical risk associated with large tumor burdens, the initial tumor burden in this study is taken up to 75% of the environmental carrying capacity 58. The initial proportion of resistant cells follows exsited evolutionary dynamical models 59. All the baseline parameter values are summarized in Table 1.
Parameter | Description | Value or range | Reference |
Total environmental carrying capacity | |||
Competition coefficient of resistant cells on sensitive cells | |||
Competition coefficient of sensitive cells on resistant cells | |||
Proliferation rate of sensitive cells | 35 | ||
Proliferation rate of resistant cells | 35 | ||
Death rate of sensitive cells | |||
Death rate of resistant cells | |||
The killing strength of therapeutic agents | 57 | ||
Maximal tolerance dose | |||
Initial tumor burden | 58 | ||
Initial proportion of resistant cells | 59 |
The treatment-pausing threshold is crucial for the efficiency of adaptive therapy
To investigate the effect of treatment-pausing thresholds on the outcomes of adaptive therapy, a simplified assumption of neutral competition () is considered. The outcomes are primarily characterized by the time to disease progression (TTP), which is typically defined as the time at which the tumor burden grows to a certain percentage greater than its initial burden35, 36, 57, 58, 59. For simplicity, this percentage is set to % in this study.
As illustrated in the example of tumor evolution under adaptive therapy (FigureΒ 2A), during the initial phase of adaptive therapy, drugs are administered to suppress tumor growth primarily through the depletion of sensitive cells. During this stage, the proportion of sensitive cells is high, which aligns with real-world cases. Following the principles of adaptive therapy, treatment is paused when the tumor burden reaches a predefined treatment-pausing threshold , conventionally set at % of the initial tumor burden, i.e., . The pause allows sensitive cells to repopulate, leading to tumor regrowth. When the tumor burden returns to the initial level , treatment resumes. The treatment-pausing threshold thus defines a cycle, from the start of treatment to its suspension. These cycles continue until resistant cells dominate the tumor burden, ultimately leading to disease progression.
To further elucidate this phenomenon, phase-plane analysis reveals distinct dynamical regimes with and without treatment (FigureΒ 2B1-B2). In untreated systems, tumor populations evolve toward coexistence equilibria, as indicated by the nullcline, which mathematically corresponds to the diagonal in FigureΒ 2B1. In the presence of treatment (FigureΒ 2B2), the long-term steady state shifts to the upper-left corner, where resistant cells dominate the tumor, and sensitive cells eventually disappear. From a dynamical perspective, treatment exerts selective pressure, pushing the tumor toward a resistant cell-dominant state. Therefore, the significance of treatment pauses lies in allowing the tumor to remain longer in a state dominated by sensitive cells, thereby delaying the transition to resistance. In this regard, increasing the treatment-pausing threshold can slow the transition of tumor composition to a resistant state, which implies it is possible to improve adaptive therapy outcomes by modulating the threshold and preventing continuous treatment from driving the tumor toward resistance via Darwinian selection 16.
Therefore, the effects of varying the treatment-pausing threshold are examined under different initial tumor states, characterized by the tumor burden () and the fraction of resistant cells () (ref. TableΒ 1). As shown in FiguresΒ 2C1-C3, altering the treatment-pausing threshold significantly impacted TTP. Notably, when the threshold is too low, meaning the tumor never reached the treatment pause threshold even during the initial cycle, adaptive therapy lacks treatment holidays and effectively becomes equivalent to MTD therapy. Consequently, as is round , TTP of AT is de facto that of MTD therapy, as shown in the leftmost segments of FiguresΒ 2C1-C3.
Importantly, for different initial tumor states, three distinct scenarios emerged when comparing adaptive therapy with MTD. The first scenario, uniform-decline (FigureΒ 2C1), occurs when adaptive therapy consistently performs worse than MTD, regardless of the threshold. This scenario typically arises in tumors with a low initial tumor burden. The second scenario, conditionally-improve (FigureΒ 2C2), occurs when adaptive therapy outperforms MTD only at specific threshold values, suggesting that careful threshold selection is required. The third scenario, uniform-improve (FigureΒ 2C3), occurs when adaptive therapy is always superior to MTD.
These findings highlight the complex relationship between the treatment-pausing threshold and TTP. Moreover, as shown in FiguresΒ 2C2-C3, the effect of the threshold is neither monotonic nor continuous, exhibiting discontinuities as the number of treatment cycles changes discretely (Figure S1). Specifically, in tumors with a low initial burden (FiguresΒ 2C1 and C2), increasing the number of treatment pauses reduces drug exposure, which in turn weakens tumor suppression because under these conditions, the initial tumor burden is low, and therefore, environmental resources remain abundant which makes the changes in the burden of sensitive cells have little impact on the growth rate of resistant cells. In contrast, for tumors with a higher initial burden (FigureΒ 2C3), where environmental resources are constrained, fluctuations in sensitive cell numbers significantly affect the growth of resistant cells. In such cases, increasing the number of treatment pauses enables sensitive cells to grow and exert stronger suppression on resistant cells, leading to a marked increase in TTP. This aligns with clinical observations in prostate cancer, where intermittent dosing preserved therapeutic sensitivity 35.
To further investigate the impact of initial tumor states on these three scenarios, we examined how different initial tumor burdens and initial resistant cell fractions influence adaptive therapy outcomes (FigureΒ 2D). When the initial tumor burden exceeds , adaptive therapy uniformly improves TTP, regardless of the initial proportion of resistant cells. However, when the initial tumor burden is below , adaptive therapy can only conditionally improve TTP, implying that careful selection of the treatment-pausing threshold is crucial. In the worst scenario, when the initial tumor burden is very low and the proportion of resistant cells is relatively high (FigureΒ 2D, top-left), adaptive therapy fails to improve TTP at any threshold, indicating primary resistance. These findings extend prior theoretical work 42 by quantifying phase-transition-like phenomena in threshold-mediated outcomes.
Furthermore, clinical implementation faces inherent uncertainty in tumor state assessment and threshold selection, implying that both the treatment-pausing threshold and the initial tumor burden and resistance fraction exhibit random variability. To address this, Monte Carlo simulations were performed to evaluate the probabilities of the three outcome scenarios under parameter variability (FigureΒ 2E). Inspired by the Norton-Simon Hypothesis 60, 61, which posits that sensitive cells typically have a greater fitness advantage than resistant cells (i.e., sensitive cells have a higher intrinsic proliferation rate than resistant cells), we further explored the effect of the relative growth rate of sensitive cells () on the probabilities of the three scenarios. By randomly varying the treatment threshold and, while keeping fixed, randomly altering and , we found that the probability of the uniform-improve scenario increases as the relative fitness of sensitive cells increases. When is small, adaptive therapy is less likely to improve TTP relative to MTD. Conversely, when is high, indicating that sensitive cells have a significant growth advantage over resistant cells, adaptive therapy is more likely to enhance TTP. These results confirm that the treatment-pausing threshold has a profound impact on adaptive therapy outcomes and patient-specific tumor characteristics, including initial tumor burden and the relative fitness of sensitive and resistant cells, significantly influence treatment outcomes. These findings align with existing research 26 and demonstrate that dependence provides quantitative criteria for patient stratification, providing insights for addressing a key challenge in empirical threshold selection 31.
Adaptive therapy is Comparison of the optimal outcomes of four therapeutic strategies
This subsection further compares the efficacy of adaptive therapy and intermittent therapy under the uniform-improve scenario, extending the previous comparison with maximum tolerated dose. As indicated in prior results, the effectiveness of different treatment strategies is influenced by treatment parameters. Therefore, to ensure a fair comparison, treatment parameters must be appropriately aligned. A reasonable approach is to compare strategies based on their best achievable outcomes (FigureΒ 3). Specifically, for intermittent therapy, the optimal treatment cycle and duration should be selected, while for adaptive therapy, the optimal threshold values should be determined.
The optimal parameter combination is determined based on different initial tumor states. Since this study does not focus on identifying these optimal parameters of intermittent therapy, computational details are omitted. However, as previous results suggest that under adaptive therapy in the conditional-improve and uniform-improve scenarios, a threshold closer to generally leads to a longer TTP (FigureΒ 3). In practical applications, a threshold near 1 would necessitate frequent initiation and cessation of treatment, requiring highly precise monitoring of tumor status, which is unrealistic. Similarly, for intermittent therapy, frequent treatment pauses and resumptions are impractical. Given these practical constraints, achieving the true optimal outcome is infeasible. Consequently, only suboptimal treatment outcomes that closely approximate the optimal results are considered as benchmarks for comparison.
For intermittent therapy, the treatment cycle period () is examined to range from to time units, with drug administration lasting from to time units () per cycle. Since this study does not focus on identifying these optimal parameters of IT, computational details are omitted. The result shows that the suboptimal period is with , simulating a regimen in which the patient undergoes both treatment and rest periods within a feasible schedule (TableΒ 1 and equationsΒ (5)). For dose-skipping adaptive therapy, the treatment-pausing threshold is set to , while for dose-modulating adaptive therapy, and are selected as the suboptimal treatment parameters (FigureΒ 3A1-A4).
In the uniform-improve scenario for adaptive therapy, a comparison of the suboptimal outcomes across the four treatment strategies (FigureΒ 3A1-A4) reveals that both adaptive therapy strategies result in a longer TTP than intermittent therapy, while intermittent therapy still outperforms MTD. Furthermore, when comparing the two adaptive therapy strategies, AT-M slightly outperforms AT-S. However, considering the operational complexity of AT-M, AT-S is primarily considered in subsequent discussions to balance efficacy with practical feasibility.
From an efficiency perspective, due to factors such as drug side effects, a higher total drug dosage is not necessarily preferable. Therefore, based on the principle of maximizing TTP with the minimum amount of drug, the contribution of unit dosage to TTP is considered a measure of treatment strategy efficiency (equationΒ (8), FigureΒ 3B1-B2):
It is observed that MTD requires the least total drug dosage (FigureΒ 3B1), but this is primarily due to its short TTP. Consequently, MTDβs efficiency is the lowest among the four treatment strategies (FigureΒ 3B2). Notably, when efficiency is used as the evaluation criterion, intermittent therapy performs the best (FigureΒ 3B2). This is because intermittent therapy, compared to adaptive therapy, has longer treatment holidays with a relatively smaller decrease in TTP, resulting in a higher contribution of unit drug dosage to TTP (FigureΒ 3B2). This suggests that intermittent therapy is a reasonable option, particularly for patients with lower drug tolerance, which is consistent with clinical practice.
Furthermore, to compare the robustness of the four treatments, the stability of TTP and treatment efficiency under uncertainties in patient physiological conditions and tumor status monitoring is examined. The impact of different initial tumor states on TTP and efficiency under the suboptimal parameters for the four strategies is analyzed (FigureΒ 3C1-C4). For different initial tumor burdens, the trends in TTP for the four treatment strategies under the suboptimal parameters are similar; TTP increases with the tumor burden, which is related to the criteria for determining tumor progression. Since progression is determined based on the initial tumor burden, a larger initial tumor burden necessitates a higher burden for progression to occur, leading to this seemingly counterintuitive result (FigureΒ 3C1). From this perspective, some researchers have proposed a βmodifying adaptive therapyβ strategy, which delays treatment until the tumor reaches a certain size 62.
Moreover, intermittent therapy shows significant variability with different initial tumor burdens (FigureΒ 3C2). When the tumor burden is small, intermittent therapy is less efficient than adaptive therapy. As the tumor burden increases, this relationship reverses. Therefore, the robustness of intermittent therapyβs efficiency is inferior to that of adaptive therapy. This suggests that when adopting intermittent therapy, a more precise assessment of the patientβs specific condition is necessary.
On the other hand, both adaptive therapy strategies and intermittent therapy show similar trends in TTP and efficiency with varying initial proportions of resistant cells (FigureΒ 3C3-C4), indicating that the optimal TTP is more influenced by cell composition when the proportion of resistant cells is very low. Notably, when the initial proportion of resistant cells is high, the TTP and efficiency of different treatment strategies are close to those of MTD. This is because, when the initial proportion of resistant cells is high, the determining factor for TTP is no longer the competition between sensitive and resistant cells but rather the proliferation of resistant cells themselves, which reduces the differences between treatment strategies.
Furthermore, similar to the previous results, considering the uncertainty in real-world scenarios, the probability of adaptive therapy being the best strategy is explored under conditions of random tumor growth rates and status monitoring. The probability that adaptive therapy results in a longer TTP than both MTD and intermittent therapy is shown in FigureΒ 3D1-D2. In real scenarios, since the tumor growth rate is unknown and tumor composition is difficult to obtain, the dependence of TTP on treatment-pausing thresholds, initial tumor burden, and initial resistant cell proportion is explored under random variations of other factors. FigureΒ 3D1 shows the dependence of the probability of adaptive therapy outperforming other treatment strategies on different thresholds ( and initial tumor burden (), with the other parameters, including the initial resistant cell proportion () and the intrinsic growth rates of tumor cells ( and ) varying randomly. FigureΒ 3D2 shows the dependence of the probability on the intial resistant fraction () and the initial tumor burden (), with other parameters, , and , varying randomly.
It can be observed that the probability of selecting adaptive therapy increases significantly when the initial tumor burden is larger. However, when the initial tumor burden is smaller, adaptive therapy does not consistently improve TTP, which aligns with previous findings. Notably, the selection of the treatment-pausing threshold and tumor composition has a weaker impact on the selection of adaptive therapy than the initial tumor burden, indicating that when detailed information about tumor components is unavailable, parameters such as the treatment threshold have less influence, with the initial tumor burden playing a key role. On the other hand, TTP is highly sensitive to the selection of the threshold, among other factors (Figures 2 and S2-S6). In other words, the selection of the optimal threshold is also dependent on various tumor states. Therefore, when other parameters are randomly varied, the impact of the threshold on TTP becomes overshadowed by fluctuations in these parameters.
These findings are consistent with clinical observations, emphasizing the importance of a precise understanding of tumor dynamics, particularly intrinsic growth rates. If such tumor information is available, treatment strategy selection can be significantly improved, highlighting the critical importance of accurate tumor characterization, especially regarding tumor dynamics.
Conditional superiority landscapes of adaptive therapy compared to other strategies
The competitive interaction between sensitive () and resistant () cells, governed by coefficients and , fundamentally determines tumor evolution dynamics under therapeutic stress 63, 64, 65, 66. Four distinct regimes emerge based on relative competition strengths: 1) weak competition (, ), 2) strong competition (, ), 3) asymmetric competition with , or 4) . Detailed analyses of and cases demonstrate that therapeutic strategies can induce resistant population extinction under specific conditions 36.
Under strong competition (, ), untreated tumors evolve toward exclusive dominance of either or , depending on initial composition (FigureΒ 4A), mathematically equivalent to equations (1)β(3) possessing two stable steady states. Continuous MTD administration drives the tumor toward a stable steady state where sensitive and resistant cells maintain a fixed ratio with unchanged tumor burden (FigureΒ 4B). In the absence of treatment, the tumorβs final state depends on its initial condition, mathematically determined by the basin of attraction in which the system starts. Thus, sensitive cell dominance at initiation leads to resistant cell elimination without treatment, requiring only therapeutic containment to prevent long-term progression. Conversely, initial resistant cell dominance under MTD results in stable tumor burden with constant cellular proportions. The initial tumor burden determines whether disease progression occurs. Even when exceeding progression thresholds, tumor burden stabilizes below carrying capacity. This property renders the βuniformly-improveβ scenario inapplicable for strong competition systems, as therapeutic outcomes primarily depend on initial tumor state rather than threshold calibration (FigureΒ 4C). That is, as long as the long-term steady state under treatment remains acceptable (FigureΒ 4B), treatment strategy selection becomes less critical for strong competition cases.
Unlike neutral or strong competition, weak competition (, ) permits coexistence equilibria regardless of initial states (FigureΒ 5A), while continuous treatment induces resistant dominance through Darwinian selection (FigureΒ 5B). Similar to neutral competition, weak competition exhibits three distinct therapeutic response scenarios (FigureΒ 5C), necessitating strategic treatment selection.
Given clinical uncertainties in weak competition systems, comprehensive parameter-space exploration proves unnecessary and inefficient. The analysis therefore focuses on measurable initial tumor burden () to evaluate adaptive therapy superiority probability under clinical uncertainty, following previous subsections (Figure 3D1-D2). Less accessible parameters including resistant cell fraction (), thresholds (), and growth rates (, ) are randomized (FigureΒ 5D0-D8). Analysis reveals enhanced adaptive therapy superiority with decreasing and increasing (FigureΒ 5D1-D8), consistent with its premise of leveraging sensitive cells to suppress resistant populations. This relationship weakens when resistant cells evade suppression.
Notably, large initial tumor burdens in strong competition systems reduce adaptive therapy selection probability (FigureΒ 5 upper-right quadrant), primarily due to: 1) bistability-mediated tumor burden homeostasis across treatments negating adaptive advantages, and 2) extended progression timelines for large burdens due to progression criteria, diminishing therapeutic differentiation. Surgical intervention may become preferable when tumor burdens approach carrying capacity. Crucially, adaptive therapy selection probability increases with initial tumor burden except at extreme values (FigureΒ 5D0), paralleling neutral competition observations (FiguresΒ 2E and 3D1-D2). These findings confirm tumor burdenβs primacy as the clinically actionable determinant while suggesting benefits from precise competition-strength measurements.
DISCUSSION
Adaptive therapy crucially leverages competitive dynamics to maintain therapeutic-sensitive cell reservoirs that suppress resistant clones, strategically delaying drug resistance evolution while minimizing tumor progression 16. By preserving sensitive cells through threshold-modulated treatment holidays, adaptive therapy sustains ecological suppression of resistant populationsβan advantage over traditional maximum tolerated dose regimens, which inadvertently accelerate resistance 29. The clinical success of adaptive therapy hinges on its paradigm-shifting principle: harnessing intra-tumor competition to transform therapy-sensitive cells into permanent controllers of resistant cell expansion 33, 67. However, the rule-of-thumb thresholds used in clinical practice may require further research and optimization, as they might not fully account for the complex, dynamic interactions between sensitive and resistant cell populations, and may vary across different tumor types and patient profiles 36, 39, 40, 68, 37, 38.
This study employs a Lotka-Volterra model to investigate how treatment-pausing thresholds influence adaptive therapy outcomes in tumors composed of drug-sensitive and resistant cells. By simulating tumor dynamics under varying thresholds, initial tumor burdens, and competition coefficients, the analysis identifies three distinct therapeutic scenarios: uniform-improve (where AT consistently outperforms MTD regardless of threshold selection), conditional-improve (AT efficacy depends on threshold calibration), and uniform-decline (MTD superiority). When considering clinical uncertainties in measurement data, generally, the greater the initial tumor burden, the more likely adaptive therapy will outperform other treatment strategies. However, in cases where the tumor burden is excessively high, the advantages of adaptive therapy become irrelevant, and alternative approaches, beyond pharmacological treatment, may be required.
The competitive dynamics between sensitive and resistant cells is also a key factor in tumor evolution under therapeutic stress. In strong competition environments, tumors tend to evolve toward a state dominated by either sensitive or resistant cells, depending on the initial composition. Continuous MTD therapy drives the tumor toward a stable state where the ratio of sensitive to resistant cells remains constant. Thus, the critical factor in treatment outcomes is not the treatment-pausing threshold but rather the impact of the initial tumor state on the final result. In weak competition environments, sensitive and resistant cells can coexist, and adaptive therapy may lead to resistant cell dominance through Darwinian selection. Treatment effectiveness is influenced by the competition coefficients, and , with lower and higher favoring adaptive therapy. Adaptive therapy prolongs TTP most effectively when sensitive cells maintain a competitive advantage and when initial tumor burden is high, enabling ecological suppression of resistant cells.
This study advances our understanding of adaptive therapy by systematically quantifying the role of treatment-pausing thresholds in modulating tumor evolution under treatment. By integrating ecological competition principles into therapeutic design, the current work provides a mechanistic basis for optimizing intermittent dosing strategies. The findings highlight that adaptive therapy is not universally superior to MTD; its success depends critically on tumor initial state and the treatment-pausing threshold selection. These insights bridge theoretical ecology and clinical oncology, emphasizing the need for dynamic, patient-specific treatment protocols rather than static dosing regimens.
The primary contribution of this work lies in establishing a quantitative relationship between treatment-pausing thresholds and progression-free survival. By categorizing outcomes into three distinct scenarios, the model offers clinicians a potential framework to stratify patients based on measurable parameters such as initial tumor burden and resistant cell fraction. Furthermore, the demonstration that strong competition can stabilize tumor burden indefinitely under specific thresholds has profound implications for managing advanced cancers. These results extend prior work 16, 42, 36, 39 by incorporating threshold-driven mechanism in adaptive therapy.
Limitations of the study
Despite these advances, the study has limitations. First, the binary classification of tumor cells overlooks phenotypic plasticity and continuous heterogeneity, which are critical drivers of therapeutic resistance 13, 14, 69. Second, model parameters such as competition coefficients (, ) are derived from population-level assumptions rather than molecular data, limiting their biological interpretability 70, 71. Third, the Lotka-Volterra framework assumes strict bistability in competitive interactions, neglecting intermediate scenarios where partial resistance or microenvironmental constraints alter dynamics 72, 73, 74. Future studies should integrate single-cell omics to parameterize competition coefficients mechanistically and explore hybrid models combining continuum and spatial interactions. Clinical validation through adaptive therapy trials incorporating real-time biomarker monitoring will be essential to translate these insights into practice.
RESOURCE AVAILABILITY
Lead contact
Requests for further information and resources should be directed to and will be fulfilled by the lead contact, Changjing Zhuge (zhuge@bjut.edu.cn).
Materials availability
This study did not generate new materials.
Data and code availability
β’
This study did not use any real-world data; all data were generated through computational simulations.
β’
All original code has been deposited on GitHub (https://github.com/zhuge-c/AT-0).
β’
Any additional information required to reanalyze the data reported in this paper is available from the lead contact upon request.
ACKNOWLEDGMENTS
This work was funded by The National Natural Science Foundation of China (NSFC) via grant 11801020.
AUTHOR CONTRIBUTIONS
Conceptualization: Changjing Zhuge, Lei Zhang; methodology: Changjing Zhuge, Haifeng Zhang; investigation, Lanfei Sun, Changjing Zhuge, Kai Kang, Xiao Wang, Haifeng Zhang; writing-βoriginal draft: Lanfei Sun, Changjing Zhuge, Lei Zhang, Haifeng Zhang; writing-βreview & editing: Leyi Zhang, Xiaoxin Wang, Kai Kang, Yanan Cai, Haifeng Zhang, Lei Zhang and Changjing Zhuge; supervision: Changjing Zhuge
DECLARATION OF INTERESTS
The authors declare no competing interests.
SUPPLEMENTAL INFORMATION INDEX
- Figure S1:
The discontinuous changes in Time to Progression (TTP) under neutral competition conditions as a result of threshold variations. The number of treatment pauses has been altered.
- Figures S2-S6:
The dependence of TTP on various parameters under neutral competition conditions.
- Figures S7-S9:
Examples comparing four distinct treatment strategies across three different scenarios under weak competition conditions, along with an analysis of TTP and efficiency index for these examples.
- Figure S10:
An example of tumor evolution dynamics under strong competition conditions, comparing AT and MTD treatments.
Supplementary figures